Three different methods for assembling a product were proposed by an industrial engineer. To investigate the number of units assembled correctly with each method, 39 employees were randomly selected and randomly assigned to the three proposed methods in such a way that each method was used by 13 workers. The number of units assembled correctly was recorded, and the analysis of variance procedure was applied to the resulting data set. The following results were obtained: SST = 13,230;SSTR = 4,550.

(a)

Set up the ANOVA table for this problem. (Round your values for MSE and F to two decimal places, and your p-value to four decimal places.)

Source
of Variation
Sum
of Squares
Degrees
of Freedom
Mean
Square
F p-value
Treatments ? ? ? ? ?
Error ? ? ?
Total ? ?

Find the value of the test statistic. (Round your answer to two decimal places.)

_____________________.

Find the p-value. (Round your answer to four decimal places.)

p-value =_________________.
Answer #1

Ans:

df(treatments)=3-1=2

df(error)=39-3=36

MS=SS/df

Source Sum Degrees Mean F p-value
of Variation of Squares of Freedom Square
Treatments 4550 2 2275 9.44 0.0005
Error 8680 36 241.11
Total 13230 38

Test statistic:

F=2275/241.11=9.44

p-value=FDIST(9.44,2,36)=0.0005